OpenStudy (anonymous):

what is the derivative of y=5x^2+4^x+6 ? Need Answer!

OpenStudy (amistre64):

theres only one of those that might be considered non-basic

OpenStudy (amistre64):

one rule that isnt covered much is: a^x = a^x ln(a) this is a general case of which: e^x = e^x ln(e) ... since ln(e) = 1 ....